The most recent financial statements for Cardinal, Inc., are shown here:

Income Statement Balance Sheet
  Sales $ 31,400   Assets $ 74,000   Debt $ 37,400
  Costs 18,450   Equity 36,600
  Taxable income $ 12,950     Total $ 74,000     Total $ 74,000
  Taxes (24%) 3,108
    Net income $ 9,842

Assets and costs are proportional to sales. Debt and equity are not. A dividend of $3,800 was paid, and the company wishes to maintain a constant payout ratio. Next year’s sales are projected to be $36,424.

What is the external financing needed?
Answer #1

Growth rate in sales=(36424-31400)/31400=16%

Dividend payout ratio=Dividend payout/Net income

=3800/9842

=0.386100386

Sales 36,424
Costs(18450*1.16) 21402
Taxable income 15022
Taxes(15022*24%) 3605.28
Net income 11416.72
Less:Dividends(0.386100386*11416.72) 4408
Addition to retained earnings 7008.72

Total assets would be=$74000*1.16=$85840

Total equity would be=$36600+Addition to retained earnings

=36600+7008.72

=$43608.72

Total assets=Total equity+Total debt

Hence external financing needed=$85840-($43608.72+37400)

=$4831.28